OpenStudy (anonymous):

find a polynomial function with these given zeros: 0,1,6

OpenStudy (anonymous):

a polynomial with zeros: 0,1,6 is given by (x-0)(x-1)(x-6) =x(x-1)(x-6) =x(x^2-x-6x+6) =x^3 - 7x^2 +6x..

OpenStudy (anonymous):

Thank you! i just wanted to make sure I got the right answer:)